A chip manufacturer tests the microprocessors produced from the production line process. The company made a statement (claim) that only 4% of its microprocessor production was damaged. To test whether the claim is true, a study was conducted by taking a sample of 577 microprocessors, of which 37 were damaged. Can the statement (claim) be rejected or accepted with a 99% confidence level?
1
Expert's answer
2020-06-15T18:23:41-0400

"Let\;\hat p=\frac{37}{577}=0.064, p_{0}=0.04,n=577\\
a) H_{0}: p \geq 0.04\\
H_{1}: p < 0.04\\
 Z_{α}=Z_{0.01}=2.326,\\
Z=\frac{\hat p -p_{0}}{\sqrt{\frac{ p_{0}(1-p_{0})}{n}}}\\
=\frac{0.064 -0.04}{\sqrt{\frac{0.04(0.96)}{577}}}=2.94\\
\text{the rejection region} :\; Z<-2.326\\
\text{The decision is to accept}:\; H_0\\
\text {This mean that the actual }\\
\text{percentage is more than or equals 0.04. }\\"